OpenStudy (anonymous):

What is the solution to x=sqrt(2x+3)

OpenStudy (abb0t):

Square both sides to get rid of the square root.

OpenStudy (anonymous):

so let me see if im getting what your saying... so square 2x and 3 right

OpenStudy (abb0t):

you have: \(\sf x^2=2x+3\). Which looks like....a polynomial that you can factor.

OpenStudy (abb0t):

Remember that \(\sf \sqrt{x} = x^\frac{1}{2}\), so when you square it, you have: \(\sf [x^\frac{1}{2}]^2\) Which is basically \(\sf x^1=x\)